OpenStudy (anonymous):

a music store is selling 2 different CDs by a local group. The group's first CD, awakening, sells for $16. their second CD, Encore, sells for $12. The stroe sells 240 of the group's CD's for a total of $3120. a. If "x" is the number of awakening CDs that were sold, write and expression for the number of encore CDs that were sold. b. write and solve an equation you can use to solve to find the number of Awakening and encore CDs sold. c. If the Awakening CDs were $18 instead of 16, would it still be possible for 240 CDs to be sold for a total of #3120? if so, how many of each were sold? justify

OpenStudy (anonymous):

A. y=240-x B. 16x+12y=3120 16x+12(240-x)=3120 4x=240 x=60 y=180 C. 18x+12(240-x)=3120 6x=240 x=40 y=200 So yes, it is possible